# Granger Causality in High-Dimensional Networks of Time Series

### v2

@@ IV Application to EEG data
    In Figure 1, we illustrate an example of one epoch from the dataset. The obtained three principal component (PC) scores are plotted over the time series data from which they were derived. This illustrates that the PC scores can capture the temporal structures/dynamics of the high-dimensional time series from which they originate and preserve the main characteristics of the time series.
    Figure 1: Isolating Nodes via Spectral Dynamic PCA method
    We compared the time-dependent statistical properties before and after applying the spectral dynamic principal component analysis (PCA) method to the channels of interest. As illustrated in Figures 2, 3 and 4, although the temporal characteristics of the series maintained, there is a noticeable trend toward stationarity in the serial correlation amplitudes, which showed signs of non-stationarity before. This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
@@
    Test results given in Table II, reveal significant differences in GC tests conducted before and after the isolation process. In several instances, the GC tests showed contrasting results after the isolation (i.e., dimension reduction), underlining the presence of spurious GC regarding 64-dimensional VAR-based GC analysis. We observed that some of the causal relationships detected in the pre-isolation (i.e., 64-dimensional VAR) were not present in the post-isolation (i.e., 6-dimensional VAR). This suggests that the influences of other channels, when not addressed properly, can lead to misleading conclusions about causality. On the other hand, estimating causal relationships in networks with many time series —using VAR models and Granger Causality tests- is quite challenging, both statistically and computationally. The main issue is the ”curse of dimensionality.” As the number of variables increases, the number of parameters we need to estimate grows rapidly, leading to a paralyzing parameter space. This can result in overfitting, higher variance in estimators, and problems like multicollinearity when variables are highly correlated. Conventional estimation methods like Ordinary Least Squares (OLS) become unfeasible in this context. Additionally, selecting the right lag order and variables is more complicated with so many variables, increasing the risk of misspecification and inflated type I errors due to multiple testing. On the computational side, high-dimensional VAR models require significant resources that burst computational time when dealing with multiple epochs and subjects. This makes the Granger Causality testing process less reliable and efficient in these high-dimensional settings.
    TABLE II: Comparison of GC Before and After Applying sDPCA
    Direction of GC GC test w.r.t 64-dim VAR (Before) GC test w.r.t 6-dim VAR (After)
    1C3 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
    1C3 $\rightarrow$ AF3 $\blacksquare$ $\square$
    1C3 $\rightarrow$ AF4 $\square$ $\blacksquare$
    1C3 $\rightarrow$ O1 $\blacksquare$ $\blacksquare$
    1C3 $\rightarrow$ O2 $\blacksquare$ $\square$
    1C4 $\rightarrow$ C3 $\square$ $\blacksquare$
    1C4 $\rightarrow$ AF3 $\square$ $\square$
    1C4 $\rightarrow$ AF4 $\square$ $\blacksquare$
    1C4 $\rightarrow$ O1 $\square$ $\square$
    1C4 $\rightarrow$ O2 $\square$ $\blacksquare$
    AF3 $\rightarrow$ AF4 $\square$ $\blacksquare$
    AF3 $\rightarrow$ C3 $\square$ $\square$
    AF3 $\rightarrow$ C4 $\square$ $\square$
    AF3 $\rightarrow$ O1 $\square$ $\square$
    AF3 $\rightarrow$ O2 $\square$ $\blacksquare$
    AF4 $\rightarrow$ AF3 $\square$ $\blacksquare$
    AF4 $\rightarrow$ C3 $\square$ $\square$
    AF4 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
    AF4 $\rightarrow$ O1 $\square$ $\square$
    AF4 $\rightarrow$ O2 $\square$ $\square$
    FO1 $\rightarrow$ O2 $\square$ $\blacksquare$
    FO1 $\rightarrow$ AF3 $\blacksquare$ $\square$
    FO1 $\rightarrow$ AF4 $\square$ $\square$
    FO1 $\rightarrow$ C3 $\blacksquare$ $\blacksquare$
    FO1 $\rightarrow$ C4 $\blacksquare$ $\square$
    FO2 $\rightarrow$ O1 $\square$ $\blacksquare$
    FO2 $\rightarrow$ AF3 $\square$ $\blacksquare$
    FO2 $\rightarrow$ AF4 $\square$ $\blacksquare$
    FO2 $\rightarrow$ C3 $\square$ $\blacksquare$
    FO2 $\rightarrow$ C4 $\square$ $\square$
    **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.      **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
    In this section, we also estimate and present the connectivities for COI in the context of data from 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks. The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the ME and MI tasks. The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels. Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
@@
    (a) Connectivity at Left Fist MI
    (b) Connectivity at Left Fist ME
    (c) Connectivity at Right Fist MI
    (d) Connectivity at Right Fist ME
    (e) Connectivity at Both Fists MI
    (f) Connectivity at Both Fists ME
    (g) Connectivity at Both Feet MI
    (h) Connectivity at Both Feet ME
    Figure 5: Connectivity at MI vs Connectivity at ME.
@@
    When comparing motor imagery and execution for both fists, the connectivity patterns showed that channel C4 had extensive Granger causality to other channels during imagery, indicating the right motor cortex’s central role in coordinating bilateral imagined movements. During execution, there was increased causality from AF3 and AF4 (pre-frontal cortices) to motor and occipital regions, suggesting greater involvement of executive functions and planning during actual movement. For both feet tasks, the connectivity was less pronounced, with fewer significant Granger causalities observed during both imagery and execution. This may imply that foot movements engage different neural pathways or that the linear assessment is insufficient to capture the complexity.
@@
    V Discussion and Conclusion
    Our study aimed to enhance the accuracy of Granger causality analysis in high-dimensional time-series networks, focusing on EEG datasets. By using the sDPCA method, we effectively isolated the channels of interest from the influence of other channels in the network. Results from both simulation studies and real EEG signal analysis demonstrate the simplicity and effectiveness of our proposed method.
    We found that the method maintained high accuracy with a moderate number of influencing nodes. However, its accuracy decreased as the number of influencing nodes increased or when the network approached super connectivity, like swarm-like neuronal connections. This highlights the sensitivity of Granger causality to external influences and emphasizes the need to accurately model such complexities in real-world data.
    Applying the proposed approach to EEG signal networks provided empirical validation of its utility. The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals. After applying sDPCA, we observed a convergence towards stationarity in the serial correlation amplitudes, bringing the data closer to the assumptions required for reliable Granger causality testing. Comparing the results before and after isolation revealed spurious causality identifications in the pre-isolation data, underlining the importance of removing extraneous influences for accurate causality assessment.
    Despite these promising findings, there are some limitations to our study. The simulation framework, while comprehensive, may not capture all nuances of real EEG data. Additionally, the test results can be significantly affected by the choice of hyperparameters, such as the number of principal components, the weights, and the number of influencing nodes. These parameters need to be carefully selected based on the specific characteristics of the datasets under study. Furthermore, our method relies on linearity-based PCA in the frequency domain, assuming that the underlying temporal dynamics can be adequately captured, which may not always hold true. Future research should consider integrating sDPCA with approaches that capture nonlinear relationships. To further improve the robustness and accuracy of causal inference in complex networks, exploring other dimensionality reduction techniques and advanced Granger causality methods would be beneficial. Moreover, the applicability of the methodology could be enhanced by including more sophisticated models of external influences and by incorporating real-time analysis capabilities.
    Overall, our proposed approach offers a practical solution for isolating and analyzing causal associations in high-dimensional time series networks. Its ability to accurately identify causal connections while mitigating the effects of confounding influences makes it a valuable tool for researchers in neuroscience and other fields dealing with complex signals. The encouraging results of the study suggest that this method can improve the reliability of causality analysis, leading to more accurate and insightful investigations of the causal dynamics in high-dimensional networks.
    VI Acknowledgments
    We acknowledge the assistance of Grammarly and DeepL AI tools for grammar checks and language enhancement during the preparation of this manuscript.

### v3

@@ 4 Application to EEG data
In Figure 3(b), we illustrate an example of one epoch from the dataset. Three sDPCA scores obtained from excluded nodes are plotted over the time series data from which they were derived. This illustrates that the small number of scores can capture the global modes and preserve temporal structures/dynamics of the time series.
Figure 3: (a)  : COI,  : Excluded nodes (b) sDPCA scores from Excluded nodes
We compared the time-dependent properties before and after applying sDPCA to the channels of interest. As illustrated in Figures 4, 5 and 6, although the temporal characteristics of autocorrelation are maintained, there is a noticeable trend toward less correlatedness at long lags, suggesting that the shared variance with other brain areas has been removed. This can be considered as an indication of isolation from influencing nodes.
@@
The remainder of this section presents estimated connectivities among COI regarding 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks. The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the ME and MI tasks. The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels. Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
@@
(a) Left Fist (MI)
(b) Right Fist (MI)
(c) Both Fists (MI)
(d) Both Feet (MI)
(e) Left Fist (ME)
(f) Right Fist (ME)
(g) Both Fists (ME)
(h) Both Feet (ME)
Figure 7: Connectivity under Motor Imagery (top row) vs. Motor Execution (bottom row). Each panel (a)–(h) shows connectivity for Left Fist, Right Fist, Both Fists, and Both Feet.
@@
When comparing motor imagery and execution for both fists, the connectivity patterns showed that channel C4 had extensive Granger causality to other channels during imagery, indicating the right motor cortex’s central role in coordinating bilateral imagined movements. During execution, there was increased causality from AF3 and AF4 (pre-frontal cortices) to motor and occipital regions, suggesting greater involvement of executive functions and planning during actual movement. For both feet tasks, the connectivity patterns involve interactions among motor, pre-frontal, and occipital channels in both imagery and execution. Imagery and execution paths share influence from C4 onto multiple channels, indicating a common mechanism connecting the right motor cortex to pre-frontal and occipital regions. Findings suggest that BF movements share overlapping causal networks in both imagery and execution, but the directionality and intensity slightly differ.
@@
5 Discussion and Conclusion
Our study aimed to enhance the usage of conventional Granger causality analysis in high-dimensional time-series networks, including application on EEG datasets. To this end, we investigated whether a two‑step strategy -sDPCA for dimensionality reduction followed by tests for Granger causality- can recover causal links between selected time series within a high-dimensional network. By projecting the activity of hundreds of background channels onto a handful of dynamic principal scores, the procedure lowers the parameter count, shortens computation time, and maintains the temporal information essential for EEG-like brain monitoring data. Simulated data confirmed that the proposed approach keeps false discoveries low and detects underlying connections if the background influence is not overwhelmingly strong. The EEG study provided a secondary line of evidence: after removing global confounding activity with sDPCA, the remaining channel pairs showed causal patterns corresponding to the known physiology of motor tasks. The results suggest that the approach captures spatial and temporal structure well enough to dissect meaningful interactions while avoiding the cost of fitting a full high‑dimensional VAR.
Several conclusions emerge. First, spectral decomposition supplies a compact set of scores that retain lag and frequency-based information under low-dimensional representation, allowing conventional F‑tests to remain valid. Second, the reduction step adds little overhead; its computational burden scales mainly with the Fourier transform used to build the spectral matrix, not the square of the channel count. Third, the isolation step mitigates the long lag serial correlation of signals, which helps to uncover connections between series more efficiently. Together, these points make the workflow practical when the goal is to focus on one or two lead-lag relations embedded in a large network.
The same findings also mark the limitations of the framework. When many channels share very strong drivers, relevant variance spreads across more dynamic components than can be kept without restoring noise, and detection power falls. Results also depend on a stable spectral density; abrupt shifts in frequency content undermine the eigen‑decomposition. Standard sDPCA relies on symmetric filters that are unusable in our context; replacing them with one‑sided causal filters prevents leakage of future information at the cost of lower frequency resolution. Finally, researchers interested in global connectivity still need sparse multivariate models. The proposed approach is not a universal replacement for full network modeling. When the scientific aim extends from a pair of nodes to the global network topology, sparse high‑dimensional modeling strategies remain necessary.
Future extensions could embed the current sDPCA setting within richer pipelines. One path is to follow the causal scores with sparse‑ or ridge‑regularised VAR estimation, while another is to expand the score vector with kernel or other nonlinear transformations so that interaction terms and threshold effects enter the GC testing. A sequential variant could let the one‑sided filter update online, allowing real‑time monitoring. The dimensionality‑reduction stage can be upgraded by forming a weighted blend of latent factors drawn from nonlinear ICA, autoencoder embeddings, and graph‑spectral filters; this composite factor set feeds into the same purge‑then‑GC workflow and equips the method to address a wider spectrum of confounding patterns.
In summary, sDPCA, followed by Granger causality testing, offers a practical compromise between full‑network modeling and bivariate analyses. It delivers reproducible inference on targeted node pairs while controlling for high‑dimensional background activity, provided that stationarity and moderate sparsity hold. These qualities make the approach useful for the toolkit for studying directed interactions in large‑scale time‑series data, such as EEG.